Materials Quantity Variance. Sweets Company produces boxes of chocolate. A standard of 2 pounds of material is expected to be used for each box produced, at a cost of $5 per pound. Sweets produced 1,000 boxes of chocolate during the month of April and used 2,200 pounds of chocolate. Calculate the materials quantity variance for the month of April.
Answer #1

Material Quantity variance = (Standard quantity-Actual quantity)*standard Price

Given information:

No of actual units produced (A)

       1,000

material required for 1000 units (1000*2 pounds) (B=A*2)

       2,000

Actual quantity used (C)

       2,200

Standard price per pound (D)

$5

Material Quantity variance = (E=(B-C)*D)

($1,000)

Ans

Material Quantity variance = 1,000 Unfavourable (2000-2000)*$5

Formula:

Material Quantity variance = (Standard quantity-Actual quantity)*standard Price